What is the rate at wich hydrogen is disappeared from the rate of formation of ammonia for the reaction H2(g)+N2(g)=2NH3(g) is 1.12 mol l-1hr-1
1
Expert's answer
2020-06-04T10:30:36-0400

N2​+3H2→2NH3​

−dtd[H2=0.3×10−4Ms−1

Rate =31​dtd[H2​]​=+21

​dtd[NH3]=dtd[NH3​]=3​dtd[H2​]=dt

d[NH3​]​=32×(0.3×10−4)

dtd[NH3​]=−0.2×10−4Ms−1